Circular No. 9120

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

V1310 SCORPII = NOVA SCORPII 2010 As first announced on CBET 2183, K. Nishiyama (Kurume, Fukuoka-

ken, Japan) and F. Kabashima (Miyaki-cho, Saga-ken, Japan) reported the discovery of a possible nova (mag 10.5) on two unfiltered CCD frames taken around Feb. 20.857 UT using a 105-mm f/4 camera lens (+ SBIG STL6303E camera); they confirmed the appearance of the new object at mag 10.3 on six unfiltered CCD frames taken around Feb. 21.832 (and mag 10.7 on Feb. 22.846) using a Meade 0.40-m reflector (+ SBIG STL1001E camera), from which they measured the following precise position: α = 17^{h}06^{m}07^{s}.53, δ = −37^{o}14′27′′.4 (equinox 2000.0). Nothing is visible at this position on two past survey frames taken on 2009 Sept. 18.425 (limiting mag 13.6) and Oct. 3.414 (limiting mag 12.3), or on a Digitized Sky Survey (DSS) plate from 1997 Apr. 2.8 (limiting red mag 18.0). Following posting on the Central Bureau’s unconfirmed-objects webpage, other observers have reported observations of the new variable. S. Kiyota (Tsukuba, Japan; remotely using a Global-rent-a-scope 32-cm reflector + SBIG ST-10XE CCD camera located at Moorook, South Australia) reports position end figures 07^{s}.51, 26′′.9 and the following magnitudes: Feb. 22.701, V = 12.03; 22.702, B = 13.09; 22.703, R_{c} = 10.70; 22.705, I_{c} = 9.65. Kiyota adds that nothing is visible at this position on a V -band Digitized Sky Survey plate from 1988 Apr. 14 (no limiting magnitude provided). E. Guido and G. Sostero also imaged this object remotely through a 0.40-m reflector from the GRAS Observatory at Moorook on Feb. 22.8, yielding position end figures 07^{s}.50, 27′′.7 and unfiltered CCD mag ∼ 10.1. W. Vollmann, Vienna, Austria, reports the following green (approximately visual) magnitudes for the variable from DSLR images taken with a 0.60-m f/3.5 reflector by H. Koberger, Vilaflor, Tenerife: Feb. 24.279, 12.3; 25.254, 12.4. Koberger adds that low-resolution spectra (grating 100 lines/mm) taken around the same times clearly show strong Hα emission.

K. Ayani, Bisei Astronomical Observatory (BAO), writes that a lowresolution spectrum (range 400–800 nm; resolution 0.5 nm at Hα) of the new variable, obtained with the BAO 1-m telescope on Feb. 23.88 UT, shows prominent Hα emission (FWHM ∼ 1800 km/s; equivalent width 70 nm) with P-Cyg absorption upon a red continuum; the absorption bottom is blueshifted by 1600 km/s with respect to the emission peak. The spectrum also shows emission lines of Hβ and Fe II (multiplet 42), which are weak and probably due to large interstellar extinction. These features suggest that the new object is a “Fe II”-type nova.

E. Kazarovets and N. N. Samus report that the GCVS designation V1310 Sco has been assigned to this nova.

2010 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green